# A Catalog of Micro Frontends Anti-patterns

### v1

@@ III Methodology
Figure 1 illustrates the process of proposing and refining the MFE anti-patterns, which we detail in this Section. Initially, we used the Systematic Literature Review (SLR) of Cerny et al. [14] as a basis for identifying microservices anti-patterns. This SLR provided a comprehensive catalog of 58 disjoint anti-patterns derived from 203 originally identified anti-patterns for microservices. Drawing upon the microservice anti-patterns identified by Cerny et al. and further explored in Taibi et al. [9], Tighilt et al. [20], Parker et al. [21] and Bogner et al. [22], coupled with an analysis of real-world problems faced by MFE development teams, we proposed a catalog of 12 MFE anti-patterns. Each anti-pattern follows the Mini-AntiPattern format (name, problem, and solution) and includes an example to illustrate the problem within an MFE context. The initial proposal of the anti-patterns can be found in our supplementary material [^1] https://github.com/nabsonp/ICSE-2025/blob/main/Initial_MFE_Anti-patterns_Catalog.pdf. We classified the anti-patterns into four categories:
@@
To evaluate the proposed anti-patterns, we conducted a survey that included both open and closed questions. The intended audience for the survey included software industry practitioners with experience in MFE development. This approach allowed the validation of different aspects of the anti-patterns by incorporating the insights of practitioners specialized in the MFE field. We also included a consent form as part of our survey instruments, allowing participants to provide their consent to participate. We assured participants that their participation was voluntary and that they could withdraw at any time. All collected data was anonymized, ensuring privacy and enabling participants to share valuable insights without concern. Our survey comprised three sections and can be found in our supplementary material [^2] https://github.com/nabsonp/ICSE-2025/blob/main/Survey_Questions.pdf.
@@ IV Results
On a brief overview of the participants’ characterization, 75% of the participants are currently involved in MFE software projects, and 75% have more than two years of experience working with MFE as well. A total of 45% of the subjects identified themselves as Fullstack Developers, 35% as Frontend Developers, 10% as Software Architects, and 5% as Mobile Developers and SE Team Leaders each. We assigned each participant an identifier ranging from P1 to P20. More than half of the participants work at large companies with over 90 software engineers and projects involving 40+ MFEs and 70+ MS, with engineering teams structured as independent units. The remaining participants work on other software companies, but in smaller projects. This diversity ensures the identified anti-patterns are applicable across different scales and organizational structures. The participants’ full characterization is available at our supplementary material [^7] https://github.com/nabsonp/ICSE-2025/blob/main/Survey_Participants_Characterization.pdf.
@@ IV-A Quantitative Analysis
We then used the Friedman test to evaluate whether there were statistically significant differences in the harmfulness scores. The Friedman test yielded a $p-value$ of $0.0000001581$ $(<0.05)$, which strongly suggests that there are statistically significant differences in harmfulness scores among the anti-patterns. Subsequently, the post-hoc Dunn test revealed that only two anti-patterns exhibited statistically significant differences compared to others: Golden Hammer differed from Hub-like Dependency, No CI/CD, Lack of Skeleton and Common Ownership; and Micro Frontend as the Goal differed from Hub-like Dependency. The implementation of the statistical tests and their results are available in our supplementary material [^8] https://github.com/nabsonp/ICSE-2025/blob/main/Statistic_tests.ipynb.
@@ IV-B Thematic Analysis
Based on the open coding quotations, the first author defined eight themes (Figure 3). Subsequently, the second author independently categorized the same quotations using the established themes. On measuring the ICR, we obtained a Cohen’s Kappa with a value of 0.84. According to Landis and Koch [32], this is considered an almost perfect score, highlighting the reliability of our coding process and the robustness of the identified themes. In the following paragraphs, we present a brief explanation of each category and its main findings. The full thematic analysis is available in our supplementary material [^9] https://github.com/nabsonp/ICSE-2025/blob/main/Thematic_Analysis.pdf.
@@ V Micro frontends anti-patterns
In this section, we present the improved version of the MFE anti-patterns, including refinements, after analyzing practitioners’ feedback. Due to the page constraint, we omitted the examples of the anti-pattern. These examples are available in our supplementary material [^10] https://github.com/nabsonp/ICSE-2025/blob/main/Improved_MFE_Anti-patterns_Catalog.pdf and at our web application [^11] https://mfe-anti-patterns.online/micro-frontends-anti-patterns/#/catalog.
@@ VIII Related work
Although some practitioners have shared their experience on anti-patterns in MFE applications on blogs and keynotes [34, 35, 36, 37], no scientific studies have been conducted to propose a catalog of MFE anti-patterns. We conducted a Rapid Review [^12] https://github.com/nabsonp/ICSE-2025/blob/main/Rapid_Review.pdf and found no studies proposing anti-patterns for MFEs. Mezzalira [6] briefly mentions that sharing any state between micro-frontends is considered an anti-pattern but does not provide an in-depth definition or exploration of this or other anti-patterns. Peltonen et al. [7] conducted a Multivocal Literature Review and stated that researchers have not yet deeply investigated MFEs, and patterns and anti-patterns have not been defined. Taibi et al. [8] presented a set of development approaches based on their experience and reported the lack of pattern and anti-pattern definitions. Moraes et al. [24] reported a case study findings, which revealed that involuntary anti-patterns may occur since they are not yet mature, generating severe negative impacts on software projects.
@@ X Data availability
The initial and improved version of the anti-patterns, including examples, survey responses, and the qualitative and quantitative analysis conducted in this research, are currently maintained as an open-source project accessible at: https://github.com/nabsonp/ICSE-2025.
@@ Acknowledgment
We thank all the practitioners who volunteered to participate in the survey for their support. The present work is the result of the Research and Development (R&D) project 001/2020, signed with the Federal University of Amazonas and FAEPI, Brazil, which has funding from Samsung, using resources from the Informatics Law for the Western Amazon (Federal Law nº 8.387/1991), and its disclosure is in accordance with article 39 of Decree No. 10.521/2020. Also supported by CAPES – Financing Code 001, CNPq process 314797/2023-8, and Amazonas State Research Support Foundation – FAPEAM – through POSGRAD 22-23.

### v4

@@ III Methodology
Figure 1 illustrates the process of proposing and refining the MFE anti-patterns, which we detail in this Section. Initially, we used the Systematic Literature Review (SLR) of Cerny et al. [14] as a basis for identifying microservices anti-patterns. This SLR provided a comprehensive catalog of 58 disjoint anti-patterns derived from 203 originally identified anti-patterns for microservices. Drawing upon the microservice anti-patterns identified by Cerny et al. and further explored in Taibi et al. [9], Tighilt et al. [20], Parker et al. [21] and Bogner et al. [22], coupled with an analysis of real-world problems faced by MFE development teams, we proposed a catalog of 12 MFE anti-patterns. Each anti-pattern follows the Mini-AntiPattern format (name, problem, and solution) and includes an example to illustrate the problem within an MFE context. The initial proposal of the anti-patterns can be found in our supplementary material (Section X). We classified the anti-patterns into four categories:
@@
To evaluate the proposed anti-patterns, we conducted a survey that included both open and closed questions. The intended audience for the survey included software industry practitioners with experience in MFE development. This approach allowed the validation of different aspects of the anti-patterns by incorporating the insights of practitioners specialized in the MFE field. We also included a consent form as part of our survey instruments, allowing participants to provide their consent to participate. We assured participants that their participation was voluntary and that they could withdraw at any time. All collected data was anonymized, ensuring privacy and enabling participants to share valuable insights without concern. Our survey comprised three sections and can be found in our supplementary material (Section X).
@@ IV Results
On a brief overview of the participants’ characterization, 75% of the participants are currently involved in MFE software projects, and 75% have more than two years of experience working with MFE as well. A total of 45% of the subjects identified themselves as Fullstack Developers, 35% as Frontend Developers, 10% as Software Architects, and 5% as Mobile Developers and SE Team Leaders each. We assigned each participant an identifier ranging from P1 to P20. More than half of the participants work at large companies with over 90 software engineers and projects involving 40+ MFEs and 70+ MS, with engineering teams structured as independent units. The remaining participants work on other software companies, but in smaller projects. This diversity ensures the identified anti-patterns are applicable across different scales and organizational structures. The participants’ full characterization is available at our supplementary material (Section X).
@@ IV-A Quantitative Analysis
We then used the Friedman test to evaluate whether there were statistically significant differences in the harmfulness scores. The Friedman test yielded a $p-value$ of $0.0000001581$ $(<0.05)$, which strongly suggests that there are statistically significant differences in harmfulness scores among the anti-patterns. Subsequently, the post-hoc Dunn test revealed that only two anti-patterns exhibited statistically significant differences compared to others: Golden Hammer differed from Hub-like Dependency, No CI/CD, Lack of Skeleton and Common Ownership; and Micro Frontend as the Goal differed from Hub-like Dependency. The implementation of the statistical tests and their results are available in our supplementary material (Section X).
@@ IV-B Thematic Analysis
Based on the open coding quotations, the first author defined eight themes (Figure 3). Subsequently, the second author independently categorized the same quotations using the established themes. On measuring the ICR, we obtained a Cohen’s Kappa with a value of 0.84. According to Landis and Koch [32], this is considered an almost perfect score, highlighting the reliability of our coding process and the robustness of the identified themes. In the following paragraphs, we present a brief explanation of each category and its main findings. The full thematic analysis is available in our supplementary material (Section X).
@@ V Micro frontends anti-patterns
In this section, we present the improved version of the MFE anti-patterns, including refinements, after analyzing practitioners’ feedback. Due to the page constraint, we omitted the examples of the anti-pattern. These examples are available in our supplementary material (Section X) and at our web application [^5] https://mfe-anti-patterns.online/micro-frontends-anti-patterns/#/catalog.
@@ VIII Related work
Although some practitioners have shared their experience on anti-patterns in MFE applications on blogs and keynotes [34, 35, 36, 37], no scientific studies have been conducted to propose a catalog of MFE anti-patterns. We conducted a Rapid Review – available at our supplementary material (Section X) – and found no studies proposing anti-patterns for MFEs. Mezzalira [6] briefly mentions that sharing any state between micro-frontends is considered an anti-pattern but does not provide an in-depth definition or exploration of this or other anti-patterns. Peltonen et al. [7] conducted a Multivocal Literature Review and stated that researchers have not yet deeply investigated MFEs, and patterns and anti-patterns have not been defined. Taibi et al. [8] presented a set of development approaches based on their experience and reported the lack of pattern and anti-pattern definitions. Moraes et al. [24] reported a case study findings, which revealed that involuntary anti-patterns may occur since they are not yet mature, generating severe negative impacts on software projects.
@@ X Data availability
The initial and improved version of the anti-patterns, including examples, survey responses, and the qualitative and quantitative analysis conducted in this research, are currently maintained as an open-source project accessible at: https://zenodo.org/records/14868084.
@@ Acknowledgment
We thank all the participants in the empirical study and USES Research Group members for their support. The present work is the result of the Research and Development (R&D) project 001/2020, signed with the Federal University of Amazonas and FAEPI, Brazil, which has funding from Samsung, using resources from the Informatics Law for the Western Amazon (Federal Law nº 8.387/1991), and its dissemination is in accordance with article 39 of Decree No. 10.521/2020. This work was carried out with the support of the Coordination for the Improvement of Higher Education Personnel - Brazil (CAPES-PROEX) – Funding Code 001; CNPq process 314797/2023-8; CNPq process 443934/2023-1; and CNPq process 445029/2024-2. Additionally, this work was partially funded by the Foundation for Research Support of the State of Amazonas – FAPEAM – through the PDPG project.